OpenStudy (whitwhit):

Harriet wants to buy items worth $248.50. She is using a store coupon for 25 percent off her purchases. She has to pay 5 percent sales tax. Calculate the total cost of the items.

Directrix (directrix):

@whitwhit If you stay in the thread, I will help you.

Directrix (directrix):

The coupon for 25% off Harriet's purchases means that Harriet has to pay a cost of 75% of the purchases. That is 75% of $248.50 or .75 * $248.50.

Directrix (directrix):

Your turn. What does this equal? .75 * $248.50

Directrix (directrix):

You can copy and paste this into the Google search box and the answer will pop up. .75 * 248.50 = Do that and post what you get. Then, there is one more step and we are finished.

OpenStudy (whitwhit):

186.37

Directrix (directrix):

186.375 which rounds up to $ 186.38. On the 186.38, she has to pay 5% sales tax. So, her final bill will be this: 186.38 * 1.05 = Multiplying by 1.05 ensures that the cost of the discounted items and the 5% tax will be included.

Directrix (directrix):

Here is your last calculation which is the answer to the question. 186.38 * 1.05 = Post what you get so we can compare answers.

OpenStudy (3mar):

@whitwhit

Directrix (directrix):

He has gone. I'll ask you what I asked him. 186.38 * 1.05 = ? That is Harriett's final bill. @3mar

OpenStudy (3mar):

It would be 195.699 I think!

OpenStudy (3mar):

is it good?

OpenStudy (3mar):

@Directrix

Directrix (directrix):

$ 195.699 which is $ 195.70 for the final bill.

Directrix (directrix):

>is it good? Good.

OpenStudy (3mar):

Thanks.